OpenStudy (anonymous):

what is the solution of the equation? 1/4b=-3

OpenStudy (anonymous):

kush has it :)

OpenStudy (anonymous):

-1/12

OpenStudy (anonymous):

\[\frac{1}{4b}=-3\] \[1=-3*4b\] \[1=-12b\] \[b=\frac{1}{-12}\]

OpenStudy (anonymous):

-0.833333333

OpenStudy (anonymous):

it said thats not it

OpenStudy (anonymous):

btw: Is it 1/(4b)=-3 or (1/4)b=-3

OpenStudy (anonymous):

Incorrect. :(

OpenStudy (anonymous):

-0.0833333333

OpenStudy (anonymous):

b = -0.0833333333

OpenStudy (anonymous):

if it's: 1/(4b)=-3 --> b=-1/12 and if it's (1/4)b=-3 --> b=-12